OpenStudy (anonymous):

an = an−1 + 3an−2, a0 = 1, a1 = 2

OpenStudy (zehanz):

What do you have to do with this sequence?

OpenStudy (zehanz):

Maybe it's clearer to write it as: \[a_n=a_{n-1}+3a_{n-2}\]\[a_0=1,a_1=2\]

OpenStudy (anonymous):

its a recursive sequence

OpenStudy (anonymous):

Yeah sorry

OpenStudy (anonymous):

im trying to figure it out because the way i keep solving it is apparently wrong by the book

OpenStudy (zehanz):

Yes, I see that, but what is your question about it? Do you need terms of this sequence?

OpenStudy (anonymous):

yes how do you solve it

OpenStudy (zehanz):

It says: If you want to know a certain number of this sequence, you need the two numbers before it. Take the last number and add 3 times the number before it to get the new one. That is why there are two start numbers given.\[a_2=a_1+3a_0\]\[a_3=a_2+3a_1\]So you can calculate every number in the sequence. I don't understand what you mean by solve it.

OpenStudy (anonymous):

Okay so if i was to do it like that then .. a2 = 2 + 3(1) = 5 right?

OpenStudy (zehanz):

OK

OpenStudy (anonymous):

and a3 = 5 + 2 = 7

OpenStudy (anonymous):

is that right or wrong because my book says im wrong

OpenStudy (zehanz):

You are, because it is 5 + 3*2 = 11

OpenStudy (anonymous):

ohh true

OpenStudy (anonymous):

sorry about that last part but the book says

OpenStudy (anonymous):

dont worry i see the mistake

OpenStudy (anonymous):

lol i never took into consideration a0 and a1 as being apart of the answer so i was so confused as to how they got 1,2,5, then 11 lol

OpenStudy (zehanz):

So a4 = 11 + 3*5 = 26 a5= 26 + 3*11 = 59 etc.

OpenStudy (anonymous):

thank you lol

OpenStudy (zehanz):

YW! a0 and a1 are needed to give you a way to "start up"..

OpenStudy (anonymous):

yea